OpenStudy (anonymous):

Four rods of equal mass m each , Ab , BC , CA , DB are placed as shown. The moment of inertia of the system about B perpendicular to the plane ABC is (AB=a)

OpenStudy (anonymous):



OpenStudy (anonymous):

@JamesJ

OpenStudy (jamesj):

Obviously break down the problem rod by rod.

OpenStudy (jamesj):

And if you're really stuck, you're just going to integrate rod by rod.

OpenStudy (anonymous):

I (BC) = ma^2/6 ryt ?

OpenStudy (shubhamsrg):

Its basically a maths ques, you just need to find rod lengths, after doing that you just have to calculate moment of inertia of each rod about B and add all

OpenStudy (anonymous):

Yup . i Have Calculated the length of rods BC = a/sqrt(2) , AC = a/sqrt(2) , DB = (sqrt2) * a / sqrt3

OpenStudy (anonymous):

I (BC) = ma^2/6 ryt ?

OpenStudy (shubhamsrg):

yep

OpenStudy (anonymous):

I (DB) = ma^2/18 ?

OpenStudy (shubhamsrg):

won' that be (2/9)ma^2

OpenStudy (anonymous):

Hw ?

OpenStudy (shubhamsrg):